OpenStudy (jojo12g):

i think D A plate has a diameter of 14 inches. What is the approximate circumference of the plate? A. 21.98 in. B. 43.96 in. C. 87.92 in. D. 153.86 in.

OpenStudy (jojo12g):

@Aveline

OpenStudy (anonymous):

i dont get any of those answers

OpenStudy (aveline):

Circumference = π*d

OpenStudy (anonymous):

i computed, with a calculator, \[14\times \pi\] and got something else

OpenStudy (fortytherapper):

Are we using 3.14 as pi, or an extended version? 3.14 works

OpenStudy (anonymous):

oh doe, i see

OpenStudy (aveline):

We are using 3.14, that's the only way it will work

OpenStudy (fortytherapper):

That's alot of food for a plate, ill buy one

OpenStudy (anonymous):

it is one of your answers, but not D

OpenStudy (fortytherapper):

Yeah, it's pi x D, as Ave said

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=14pi

OpenStudy (anonymous):

Try the answers A and B. Look at those two.